# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v2

@@ Abstract
Studies of the dynamics of nonlinear recurrent neural networks often assume independent and identically distributed couplings, but large-scale connectomics data indicate that biological neural circuits exhibit markedly different connectivity properties. These include rapidly decaying singular-value spectra and structured singular-vector overlaps. Here, we develop a theory to analyze how these forms of structure shape high-dimensional collective activity in nonlinear recurrent neural networks. We first introduce the random-mode model, a random-matrix ensemble related to the singular-value decomposition that enables control over the spectrum and right-left mode overlaps. Then, using a novel path-integral calculation, we derive analytic expressions that reveal how connectivity structure affects features of collective dynamics: the dimension of activity, which quantifies the number of high-variance collective-activity fluctuations, and the temporal correlations that characterize the timescales of these fluctuations. We show that connectivity structure can be invisible in single-neuron activities while dramatically shaping collective activity. Furthermore, despite the nonlinear, high-dimensional nature of these networks, the dimension of activity depends on just two connectivity parameters—the variance of the couplings and the effective rank of the coupling matrix, which quantifies the number of dominant rank-one connectivity components. We contrast the effects of single-neuron heterogeneity and low-dimensional connectivity, making predictions about how $z$-scoring data affects the dimension of activity. Finally, we demonstrate the presence of structured overlaps between left and right modes in the Drosophila connectome, incorporate them into the theory, and show how they further shape collective dynamics.
@@ I Introduction
The collective activity of a high-dimensional nonlinear system is determined by the structure of the interactions between its elements, but the mapping from structure to activity is generally only analytically tractable for limited forms of structure. In systems neuroscience, structure in synaptic connectivity controls the function of a neural system by determining what patterns of activity are produced by populations of neurons. This paper studies this mapping in the context of nonlinear recurrent neural networks, used widely in neuroscience as models of neural-circuit dynamics [1, 2, 3] and in machine learning as systems for sequence processing [4, 5, 6, 7]. While these models omit many features of actual neural circuits, they capture several of their fundamental features, including their large scale, nonlinear units, and recurrent interactions. Functionally, a key property of such systems is that, like real neural circuits, they can generate rich time-varying activity in the absence of external input [8, 9].
@@
Theoretical studies of nonlinear recurrent neural networks, including the seminal work of Sompolinsky et al. [10], often focus on the case of independent and identically distributed (i.i.d.) couplings, or simple variants thereof [11]. Such disordered networks produce high-dimensional chaotic activity, reminiscent of asynchronous cortical activity observed in vivo [10]. Initializing recurrent neural networks so that they exhibit chaotic dynamics has also been shown to facilitate subsequent learning [12, 13, 14, 15, 16].
@@
Experimental access to large-scale connectivity structure creates an opportunity and imperative to develop theories that explain models beyond those with i.i.d. couplings. Recently, large-scale synaptic reconstructions of neural circuits, or “connectomes,” have become available in multiple species, including the full brain of Drosophila melanogaster [17] and increasingly large portions of mammalian circuits. These large-scale connectomes reveal that neural circuits deviate substantially from the i.i.d. assumption, as previously demonstrated via the presence of small-scale motifs [18]. In particular, in accordance with real-world networks more generally [19], neural circuits appear to exhibit low-rank structure in which the connectivity is well-described by fewer rank-one components than the network size (we demonstrate this explicitly for the Drosophila connectome in Sec. II.1). In neural-network models, recurrent or otherwise, such structure emerges naturally in several contexts, including when connections depend on distance in physical or feature space [20, 21, 22], or when networks are trained on tasks with low-dimensional structure [23, 24, 25].
@@
In parallel with the development of large-scale connectivity maps, recent advances in recording technologies including silicon probes [26] and large-scale calcium imaging [27] now enable the activities of many neurons to be monitored simultaneously, with the latest datasets including hundreds of thousands of neurons. This technological progress has enabled researchers to characterize collective properties of neural activity rather than focusing solely on single-neuron responses. These collective properties capture how neurons interact and reveal distributed computational processes that are only visible with high-yield recording technologies [28]. Such population-level features are typically analyzed using dimensionality reduction techniques like principal component analysis [29].
@@
Relating connectivity and collective-activity structure analytically. We currently lack analytical tools to predict relationships between collective activity described by large-scale recordings, and connectivity structure described by connectomics or other datasets. Dynamical mean-field theory (DMFT) is widely used to analytically characterize activity in large, nonlinear recurrent neural networks. Although recent advances in DMFT have extended these techniques to describe collective features in networks producing high-dimensional chaotic activity [30], these calculations have been limited to i.i.d. connectivity or connectivity with simple correlations between reciprocal couplings. Extending such calculations to more complicated connectivity structures rapidly becomes unwieldy. Consequently, such analyses have not been performed for networks whose connections are constrained by the statistics of connectomic or other datasets, limiting our ability to leverage these datasets for theoretical insight.
@@
One tractable approach has been to study networks with very low-dimensional connectivity and thus activity (specifically, in the limit where the rank of the coupling matrix remains finite while the network becomes large; see Sec. II.3). However, connectomes, like that of Drosophila considered in Sec. II.1, do not support this finite-rank assumption. Furthermore, while neural activity recorded in experiments is often low-dimensional, this may be inherited from the low dimensionality of experimental tasks; in more complicated tasks or spontaneous states, activity is typically higher-dimensional [31, 27].
@@
To address these limitations and challenges, this paper makes two contributions:
@@
    A tractable model of neural-network coupling matrices that we call the random-mode model. Motivated by the Drosophila connectome considered in Sec. II.1, we introduce the random-mode model in Sec. II.2, in which couplings are generated as a sum of rank-one outer products of random left and right modes with component-specific strengths. Unlike well studied low-rank recurrent neural network models, we scale the rank with the network size (Sec. II.3). We show in Sec. II.4 that parameterizing these strengths provides control over the singular value spectrum of the coupling matrix.
@@
    A new path-integral calculation of a specific four-point function of network activity, originally defined by Clark et al. [30], that captures features of collective activity including its dimensionality. We review these and other summary statistics in Sec. III. In Sec. IV A–C, we describe the calculation, based on fluctuations around the saddle point in a path integral, and apply it to the random-mode model. In Sec. IV.4, we use the solution to study how features of the coupling matrix control collective activity.
Additionally, in Sec. V, we analyze notable limiting cases of the random-mode model. In Sec. VI, we study a generalization of the model featuring heterogeneity among single-neuron properties and contrast these effects with those of low-rank structure in connectivity. Finally, in Sec. VII, we incorporate structured overlaps between left and right modes into the model, demonstrating their presence in the Drosophila connectome and showing how they further shape collective dynamics.
@@ II.1 Spectral structure of neural-circuit connectivity
Figure 1: Analysis of Drosophila central brain connectome. (a) Volume of fly brain for which reconstruction was performed (blue; reproduced from [17]). (b) Visualization of the full normalized coupling matrix. (c) Singular value spectra of the normalized coupling matrix (red) and an i.i.d. random matrix (gray). The fly connectome exhibits a smooth spectrum that decays quickly, corresponding to a reduced participation ratio. $N=18028$ neurons.
Let $\bm{J}$ denote an $N\times N$ weighted, directed coupling matrix among $N$ neurons. Many theoretical studies have examined single-neuron [10] and collective [30] properties of activity in nonlinear recurrent neural networks with i.i.d. $J_{ij}$. However, the coupling matrices of real-world networks, including neural circuits, often exhibit approximate low-rank structure that deviates dramatically from the i.i.d. assumption. To illustrate this, we analyzed a central-brain connectome of the fruit fly Drosophila melanogaster using the singular value decomposition (SVD), which decomposes any matrix as a sum of rank-one components, with each component’s strength given by the corresponding singular value (Fig. 1a; details in Appendix B). Specifically, the SVD decomposes $\bm{J}$ in the form
@@
The ranked singular values of the Drosophila connectome decay much more rapidly than those of an i.i.d. matrix of the same size, which follow a universal distribution independent of single-element statistics (Fig. 1c; similar analysis appears in Thibeault et al. [19]). We quantify this rapid decay using the participation ratio of the squared singular-value spectrum, $\text{PR}^{S}=(\sum_{a=1}^{M}S_{a}^{2})^{2}/\sum_{a=1}^{M}S_{a}^{4}$. The connectome’s participation ratio is 0.18, substantially lower than the 0.5 value characteristic of i.i.d. matrices, confirming that this biological network exhibits much more concentrated low-rank structure than would arise from random and independent couplings.
@@ II.2 Definition of the random-mode model
Both the singular-value spectrum and, as we will show later, the overlaps of the left and right singular vectors of the Drosophila connectome deviate from the predictions of an i.i.d. coupling matrix. We therefore introduce the random-mode model, a generative model for coupling matrices that allows us to capture these properties within a random-matrix ensemble.
The random-mode model has the same mathematical form as an SVD, but is a statistical generative process for $\bm{J}$ rather than a linear algebraic decomposition of it. The model generates a coupling matrix of the form
@@
where, for each $a=1,\ldots,M$, $D_{a}>0$ are the component strengths; $\bm{r}_{a}$ and $\bm{\ell}_{a}$ are $N$-dimensional left and right modes, respectively; and $M$ is the number of components, where $\text{rank}=\text{min}(N,M)$. We now describe the properties of these quantities.
Component strengths. The component strengths $D_{a}$ are analogous to the singular values $S_{a}$ of $\bm{J}$. We model these quantities deterministically with the requirement that they are defined for arbitrarily large $M$ and that empirical averages $M^{-1}\sum_{a=1}^{M}f(D_{a})$ converge to limiting values, denoted $\left\langle f(D)\right\rangle_{D}$, as $M\to\infty$.
Mode vectors. The left and right modes $\bm{\ell}_{a}$ and $\bm{r}_{a}$ are analogous to the left and right singular vectors $\bm{u}_{a}$ and $\bm{v}_{a}$ of $\bm{J}$. We assume they are sampled i.i.d. across the neuron index $i$:
@@
This yields orthonormality of mode vectors in expectation:
@@
in analogy to singular vectors for which orthonormality is deterministic  [^1] This random structure, rather than exact, deterministic orthonormality, offers analytical tractability by eliminating the need to enforce orthonormality—a complicated global constraint. But see [82], who handle random orthonormal matrices in a regression setting.. Individual mode components are $\mathcal{O}(1/\sqrt{N})$. Note that we have not constrained second moments that mix left and right mode components.
@@ II.3 Rank scaling and comparison to low-rank recurrent neural networks
A crucial modeling choice is how the number of modes, $M$, scales relative to the network size, $N$. The two main possibilities are $M=\mathcal{O}(N)$ or $M=\mathcal{O}(1)$, corresponding to extensive and intensive rank scaling, respectively. The random-mode model is characterized by extensive rank scaling, where both $M$ and $N$ approach infinity while maintaining a fixed ratio:
@@
By choosing $\alpha$ to be small, or by choosing the spectrum of ranked component strengths to decay rapidly, we can construct connectivity that is low-rank, or well-described by a small number of rank-one components, compared to the number of neurons.
The alternative approach of intensive rank scaling, where $M$ remains finite while $N\to\infty$, leads to what has been termed “low-rank recurrent neural networks” in the literature [33, 34, 35, 36]. In some instances, the intensive-rank component is added to an i.i.d. matrix modeling unstructured “background” connectivity [33, 34]. Note that in this paper, when we use “low-rank” or “low-dimensional,” we refer to quantities that are small as a fraction of $N$, even if they are extensive; this differs from some prior works (such as those on low-rank recurrent neural networks) where these terms refer to intensive quantities.
The distinction between extensive and intensive rank scaling has important implications for specifying the distribution over the $2M$ mode components $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$. For finite $M$, one can directly specify this distribution. At large $N$, such a distribution encodes the relative geometry of the left and right modes, since the inner products between modes converge to ($N$ times) the second moments of $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$, with negligible fluctuations. A common choice is a $2M$-dimensional multivariate Gaussian distribution. An important insight from prior work on the intensive-rank case is that these overlaps are crucial for shaping the dynamics and computations implemented by the network, as overlaps between connectivity modes dictate how activity modes interact [33, 34, 35, 36].
Our goal is to go beyond the intensive-rank setup, which produces correspondingly low-dimensional activity, to model more sophisticated, higher-dimensional dynamics. However, specifying the full distribution $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ in the random-mode model, where $M\to\infty$, is more complicated. First, it must be well-defined for all $M$ to allow taking the limit. More fundamentally, we would like to use the random-mode model to construct generative models that capture certain essential spectral features of coupling matrices, over which we have systematic control. Using a joint Gaussian distribution would require specifying $M^{2}$ parameters which, since $M=\mathcal{O}(N)$, is comparable to defining the full $N\times N$ coupling matrix.
@@
where $P(\ell_{a},r_{a})$ must be specified. Moreover, until Sec. VII, we further specialize to the case where the left and right modes are fully independent: $P(\ell_{a},r_{a})=P(\ell_{a})P(r_{a})$, with each marginal factor $P(\ell_{a})$ or $P(r_{a})$ being a univariate Gaussian with mean zero and variance $1/N$  [^2] The Gaussian assumption for individual components is not crucial—our results depend only on first and second moments for large $N$ due to the central limit theorem; higher-order cumulants (scaling appropriately with $1/N$) are subdominant in the path integral saddle-point approximation.. In this case, all $2MN$ components specifying the mode vectors are i.i.d. with mean zero and variance $1/N$; we refer to this case as “i.i.d. modes.” In Sec. VII we use a non-factorized form of $P(\ell_{a},r_{a})$ to specify correlations between $\ell_{a}$ and $r_{a}$, introducing $M$ deterministic parameters (in addition to the $M$ component strengths) to specify the left–right correlation for each $a$.
II.4 Relationship to singular value decomposition
In the SVD, the left and right singular vectors are orthonormal: $\bm{u}_{a}^{T}\bm{u}_{b}=\bm{v}_{a}^{T}\bm{v}_{b}=\delta_{ab}$. In contrast, the mode vectors are orthonormal only in expectation (Eq. 6); once sampled, there are random $\mathcal{O}(1/\sqrt{N})$ overlaps:

### v3

@@ Abstract
Studies of the dynamics of nonlinear recurrent neural networks often assume independent and identically distributed couplings, but large-scale connectomics data indicate that biological neural circuits exhibit markedly different connectivity properties. These include rapidly decaying singular-value spectra and structured singular-vector overlaps. Here, we develop a theory to analyze how these forms of structure shape high-dimensional collective activity in nonlinear recurrent neural networks. We first introduce the random-mode model, a random-matrix ensemble related to the singular-value decomposition that enables control over the spectrum and right-left mode overlaps. Then, using a novel path-integral calculation, we derive analytical expressions that reveal how connectivity structure affects features of collective dynamics: the dimension of activity, which quantifies the number of high-variance collective-activity fluctuations, and the temporal correlations that characterize the timescales of these fluctuations. We show that connectivity structure can be invisible in single-neuron activities while dramatically shaping collective activity. Furthermore, despite the nonlinear, high-dimensional nature of these networks, the dimension of activity depends on just two connectivity parameters—the variance of the couplings and the effective rank of the coupling matrix, which quantifies the number of dominant rank-one connectivity components. We contrast the effects of single-neuron heterogeneity and low dimensional connectivity, making predictions about how z-scoring data affects the dimension of activity. Finally, we demonstrate the presence of structured overlaps between left and right modes in the Drosophila connectome, incorporate them into the theory, and show how they further shape collective dynamics.
@@ I Introduction
The collective activity of a high-dimensional nonlinear system is determined by the structure of the interactions between its elements, but the mapping from structure to activity is generally analytically tractable only for limited forms of structure. In systems neuroscience, structure in synaptic connectivity controls the function of a neural system by determining what patterns of activity are produced by populations of neurons. This paper studies this mapping in the context of nonlinear recurrent neural networks, used widely in neuroscience as models of neural-circuit dynamics [1, 2, 3] and in machine learning as systems for sequence processing [4, 5, 6, 7]. While these models omit many aspects of actual neural circuits, they capture several of their fundamental features, including their large scale, nonlinear units, and recurrent interactions. Functionally, a key property of such systems is that, like real neural circuits, they can generate rich time-varying activity in the absence of external input [8, 9].
@@
Theoretical studies of nonlinear recurrent neural networks, including the pioneering work of Sompolinsky et al. (Ref. [10]), often focus on the case of independent and identically distributed (i.i.d.) couplings, or simple variants thereof [11]. Such disordered networks produce high-dimensional chaotic activity, reminiscent of asynchronous cortical activity observed in vivo [10]. Initializing recurrent neural networks so that they exhibit chaotic dynamics has also been shown to facilitate subsequent learning [12, 13, 14, 15, 16].
@@
Experimental access to large-scale connectivity structure creates an opportunity and imperative to develop theories that explain models beyond those with i.i.d. couplings. Recently, large-scale synaptic reconstructions of neural circuits, or “connectomes,” have become available in multiple species, including the full brain of Drosophila melanogaster [17] and increasingly large portions of mammalian circuits [18, 19, 20, 21]. These large-scale connectomes reveal that neural circuits deviate substantially from the i.i.d. assumption, as previously demonstrated via the presence of small-scale motifs [22]. In particular, in accordance with real-world networks more generally [23], neural circuits appear to exhibit low-rank structure in which the connectivity is well described by fewer rank-one components than the network size (we demonstrate this explicitly for the Drosophila connectome in Sec. II.1). In neural-network models, recurrent or otherwise, such structure emerges naturally in several contexts, including when connections depend on distance in physical or feature spaces [24, 25, 26], or when networks are trained on tasks with low dimensional structure [27, 28, 29].
@@
In parallel with the development of large-scale connectivity maps, recent advances in recording technologies, including silicon probes [30] and large-scale calcium imaging [31, 32], now enable the activities of many neurons to be monitored simultaneously, with the latest datasets including hundreds of thousands of neurons or more [33]. This technological progress has enabled researchers to characterize collective properties of neural activity rather than focusing solely on single-neuron responses. These collective properties capture how neurons interact and reveal distributed computational processes that are visible only with high-yield recording technologies [34]. Such population-level features are typically analyzed using dimensionality reduction techniques like principal components analysis [35, 36, 37, 38].
I.1 Relating connectivity and collective-activity structure analytically.
@@
We currently lack analytical tools to predict relationships between connectivity structure, described by connectomics or other datasets, and collective activity, described by large-scale neural recordings. Dynamical mean-field theory (DMFT) is a theoretical tool from statistical physics widely used to analytically characterize activity in large nonlinear recurrent neural networks. Although recent advances in DMFT have extended these techniques to describe collective features in networks producing high-dimensional chaotic activity [39], these calculations have been limited to i.i.d. connectivity, or connectivity with simple correlations between reciprocal couplings. Extending such calculations to more complicated connectivity structures rapidly becomes unwieldy. Consequently, such analyses have not been performed for networks whose connections are constrained by the statistics of connectomic or other datasets, limiting our ability to leverage these datasets for theoretical insight.
@@
One tractable approach has been to study networks with very low dimensional connectivity and thus activity (specifically, in the limit where the rank of the coupling matrix remains finite while the network becomes large; see Sec. II.3). However, connectomes, like that of Drosophila considered in Sec. II.1, do not support this finite-rank assumption. Furthermore, while neural activity recorded in experiments is often low dimensional, this may be inherited from the low dimensionality of experimental tasks; in more complicated tasks or spontaneous states, activity is typically higher dimensional [31, 32, 33].
@@
To address these limitations and challenges, this paper makes two contributions.
@@
    A tractable model of neural-network coupling matrices that we call the random-mode model. Motivated by the Drosophila connectome considered in Sec. II.1, we introduce the random-mode model in Sec. II.2, in which the couplings are generated as a sum of rank-one outer products of random left and right modes with component-specific strengths. Unlike well-studied low-rank recurrent neural-network models, we scale the rank with the network size (Sec. II.3). We show in Sec. II.4 that parametrizing these strengths provides control over the singular-value spectrum of the coupling matrix.
@@
    A new path-integral calculation of a specific four-point function of network activity, originally defined by Clark et al. (Ref. [39]), that captures features of collective activity including its dimensionality. We review these and other summary statistics in Sec. III. In Secs. IV.1IV.3, we describe the calculation, based on fluctuations around the saddle point in a path integral, and apply it to the random-mode model. In Sec. IV.4, we use the solution to study how features of the coupling matrix control collective activity.
Additionally, in Sec. V, we analyze notable limiting cases of the random-mode model. In Sec. VI, we study a generalization of the model featuring heterogeneity among single-neuron properties and contrast these effects with those of low-rank structure in connectivity. Finally, in Sec. VII, we incorporate structured overlaps between the left and right modes into the model, demonstrating their presence in the Drosophila connectome and showing how they further shape collective dynamics.
@@ II.1 Spectral structure of neural-circuit connectivity
Figure 1: Analysis of Drosophila central-brain connectome. (a) Volume of fly brain for which reconstruction was performed (blue; reproduced from [17]). (b) Normalized coupling matrix (elements summed within $10\times 10$ blocks to aid visualization). (c) Singular-value spectra of the normalized coupling matrix (red) and an i.i.d. random matrix (gray). The fly connectome exhibits a smooth spectrum that decays quickly, corresponding to a reduced participation ratio. $N=18028$ neurons.
Let $\bm{J}$ denote an $N\times N$ weighted, directed coupling matrix among $N$ neurons. Many theoretical studies have examined single-neuron [10] and collective [39] properties of activity in nonlinear recurrent neural networks with i.i.d. $J_{ij}$. However, the coupling matrices of real-world networks, including neural circuits, often exhibit approximate low-rank structure that deviates dramatically from the i.i.d. assumption. To illustrate this, we analyzed a central-brain connectome of the fruit fly Drosophila melanogaster using the singular-value decomposition (SVD), which decomposes any matrix as a sum of rank-one components, with each component’s strength given by the corresponding singular value [Figs. 1(a) and 1(b); details in Appendix B]. Specifically, the SVD decomposes $\bm{J}$ in the form
@@
The ranked singular values of the Drosophila connectome decay much more rapidly than those of an i.i.d. matrix of the same size, which follow a universal distribution independent of single-element statistics [Fig. 1(c); similar analysis appears in Thibeault et al. (Ref. [23])]. We quantify this rapid decay using the participation ratio (PR) of the squared singular-value spectrum, $\text{PR}^{S}=(\sum_{a=1}^{M}S_{a}^{2})^{2}/\sum_{a=1}^{M}S_{a}^{4}$. The connectome’s participation ratio is 0.18, substantially lower than the value 0.5 characteristic of i.i.d. matrices, confirming that this biological network exhibits a much more concentrated low-rank structure than would arise from i.i.d. couplings.
@@ II.2 Definition of the random-mode model
Both the singular-value spectrum and, as we will show later, the overlaps of the left and right singular vectors of the Drosophila connectome deviate from the predictions of an i.i.d. coupling matrix. We therefore introduce the random-mode model, a generative model for coupling matrices that allows us to capture these properties within a random-matrix ensemble (Fig. 2).
The random-mode model has the same mathematical form as an SVD, but is a statistical generative process for $\bm{J}$ rather than a matrix factorization. The model generates a coupling matrix of the form
@@
where, for each $a=1,\ldots,M$, $D_{a}>0$ are the component strengths; $\bm{r}_{a}$ and $\bm{\ell}_{a}$ are $N$-dimensional right and left modes, respectively; and $M$ is the number of components, where $\text{rank}\leq\text{min}(N,M)$. We now describe the properties of these quantities. See Appendix A for a glossary of terms.
II.2.1 Component strengths.
The component strengths $D_{a}$ are analogous to the singular values $S_{a}$ of $\bm{J}$. We model these quantities deterministically with the requirement that they are defined for arbitrarily large $M$ and that empirical averages $M^{-1}\sum_{a=1}^{M}f(D_{a})$ converge to limiting values, denoted $\left\langle f(D)\right\rangle_{D}$, as $M\to\infty$.
II.2.2 Left and right modes.
The left and right modes $\bm{\ell}_{a}$ and $\bm{r}_{a}$ are analogous to the left and right singular vectors $\bm{u}_{a}$ and $\bm{v}_{a}$ of $\bm{J}$. We assume they are sampled i.i.d. across the neuron index $i$:
@@
This yields orthonormality of modes in expectation,
@@
in analogy to singular vectors, for which orthonormality holds exactly  [^1] This random structure, rather than exact, deterministic orthonormality, offers analytical tractability by eliminating the need to enforce orthonormality—a complicated global constraint; but see also Ref. [41], which handles random orthonormal matrices in a regression setting.. Individual mode components are $\mathcal{O}(1/\sqrt{N})$. Note that we have not constrained second moments that mix left and right mode components.
@@ II.3 Rank scaling and comparison to low-rank recurrent neural networks
A crucial modeling choice is how the number of modes $M$ scales relative to the network size $N$. The two main possibilities are $M=\mathcal{O}(N)$ or $M=\mathcal{O}(1)$, corresponding to extensive and intensive rank scaling, respectively. The random-mode model is characterized by extensive rank scaling, where both $M$ and $N$ approach infinity while maintaining a fixed ratio:
@@
By choosing $\alpha$ to be small, or by choosing the spectrum of ranked component strengths to decay rapidly, we can construct connectivity that is low rank or well described by a small number of rank-one components, compared to the number of neurons.
The alternative approach of intensive rank scaling, where $M$ remains finite while $N\to\infty$, leads to what has been termed “low-rank recurrent neural networks” in the literature [42, 43, 44, 45]. In some instances, the intensive-rank component is added to an i.i.d. matrix modeling unstructured “background” connectivity [42, 43]. Note that, in this paper, when we use “low rank” or “low dimensional,” we refer to quantities that are small as a fraction of $N$, even if they are extensive; this differs from some prior works (such as those on low-rank recurrent neural networks) where these terms refer to intensive quantities.
The distinction between extensive- and intensive-rank scaling has important implications for specifying the distribution over the $2M$ mode components $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$. For finite $M$, one can directly specify this distribution. At large $N$, such a distribution encodes the relative geometry of the left and right modes, since the inner products between modes converge to ($N$ times) the second moments of $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$, with negligible fluctuations. A common choice is a $2M$-dimensional multivariate Gaussian distribution. An important insight from prior work on the intensive-rank case is that these overlaps are crucial for shaping the dynamics and computations implemented by the network, as overlaps between connectivity modes dictate how activity modes interact [42, 43, 44, 45].
Our goal is to go beyond the intensive-rank setup, which produces correspondingly low dimensional activity, to model more sophisticated, higher dimensional dynamics. However, specifying the full distribution $P(\{\ell_{a},r_{a}\}_{a=1}^{M})$ in the random-mode model, where $M\to\infty$, is more complicated. First, it must be well defined for all $M$ to allow taking the limit. More fundamentally, we would like to use the random-mode model to construct generative models that capture certain essential spectral features of coupling matrices, over which we have systematic control. Using a joint Gaussian distribution would require specifying $M^{2}$ parameters which, since $M=\mathcal{O}(N)$, is comparable to defining the full $N\times N$ coupling matrix.
@@
where $P(\ell_{a},r_{a})$ must be specified. Moreover, until Sec. VII, we further specialize to the case where the left and right modes are fully independent: $P(\ell_{a},r_{a})=P(\ell_{a})P(r_{a})$, with each marginal factor $P(\ell_{a})$ or $P(r_{a})$ being a univariate Gaussian with mean zero and variance $1/N$  [^2] The Gaussian assumption for individual components is not crucial—our results depend on only first and second moments for large $N$ due to the central limit theorem; higher-order cumulants (scaling appropriately with $1/N$) are subdominant in the path-integral saddle-point approximation.. In this case, all $2MN$ components specifying the modes are i.i.d. with mean zero and variance $1/N$; we refer to this case as “i.i.d. modes.” In Sec. VII we use a nonfactorized form of $P(\ell_{a},r_{a})$ to specify correlations between $\ell_{a}$ and $r_{a}$, introducing $M$ deterministic parameters (in addition to the $M$ component strengths) to specify the left-right correlation for each $a$.
II.4 Relationship to singular-value decomposition
In the SVD, the left and right singular vectors are orthonormal: $\bm{u}_{a}^{T}\bm{u}_{b}=\bm{v}_{a}^{T}\bm{v}_{b}=\delta_{ab}$. In contrast, the left and right modes are orthonormal only in expectation [Eq. (6)]; once sampled, there are random $\mathcal{O}(1/\sqrt{N})$ overlaps,